# The Unbounded Denominators Conjecture

### v1

@@ 1.1.4. The dimension bound can be leveraged with growing level NN
The meaning of the requisite inflation rate is clarified in § 4, with Proposition 4.2.5 and Remark 4.2.8. It turns out that the logarithmically inflated holonomy rank (dimension) bound by $O(N^{3}\log{N})$ is sufficient for the desired proof by contradiction (but an $O(N^{3+1/\log{\log{N}}})$ or worse form of bound would not suffice); and this is what we ultimately prove. Getting to this degree of precision creates however some additional challenges. A straightforward elaboration of André’s original argument in [And89, Criterium VIII 1.6], that involves a $\sup_{|z|=r}\log{|\varphi|}$ growth term in place of our mean (integrated) growth term $m(r,\varphi)$ in (2.0.4), leads quite easily to an $O(N^{5})$ dimension bound; and by a further work explicitly with the cusps of the Fuchsian uniformization $D(0,1)/\Gamma_{N}\cong\mathbf{C}\smallsetminus\mu_{N}$, and an appropriate Riemann map precomposition, it is possible to further reduce that down to an $O(N^{4})$. See Remark 5.5.8. This does not suffice to conclude the proof. Going further requires an intrinsic improvement into André’s dimension bound itself: the reduction of the supremum term to the integrated term in the numerator of (2.0.4). We achieve this in § 2.2 by an asymptotic equidistribution idea (all under $d\to\infty$), of a familiar potential-theoretic flavor similar to the proof of Bilu’s equidistribution theorem [Bil97]. The resulting potential-theoretic connection is in the cross-variables asymptotic aspect and different than the well-established link (see [Bos99, BCL09, Bos04]) of arithmetic algebraization to adelic potential theory.
@@ 2. The arithmetic holonomicity theorem
Our proof relies on the following dimension bound which is an extension of André’s arithmetic algebraicity criterion [And04, Théorème 5.4.3]. We state and prove our result here in a particular case suited to our needs. Firstly we introduce an algebra of holonomic power series with integral coefficients and restricted singularities.
Definition 2.0.1.
For $U\subset\mathbf{C}$ an open subset, $R\subset\mathbf{C}$ a subring, and $x(t)\in t\mathbf{Q}\llbracket t\rrbracket$ a formal power series, we define $\mathcal{H}(U,x(t),R)$ to be the ring of formal power series $f(x)\in(R\otimes_{\mathbf{Z}}\mathbf{Q})\llbracket x\rrbracket$ whose $t$-expansion $f(x(t))\in R\llbracket t\rrbracket$, and such that there exists a nonzero linear differential operator $L$ over $\mathbf{C}(x)$, without singularities on $U$, with $L(f)=0$.
For $x(t)=t$, we simply denote this ring by $\mathcal{H}(U,R)$.
Throughout our paper, we will use the notation
@@
Our holonomy bound is the following. The novel point is an equidistribution argument, of a potential-theoretic flavor, for achieving the integrated term instead of a supremum term in the bound (2.0.4) below. This is critical for our proof of the unbounded denominators conjecture to have the growth term in (2.0.4) expressed as a Nevanlinna mean proximity function (or a characteristic function).
@@
Let $0\in U\subset\mathbf{C}$ be an open subset containing the origin. If the uniformization radius of the pointed Riemann surface $(U,0)$ is strictly greater than $1$, then the algebra $\mathcal{H}(U,\mathbf{Z})\otimes\mathbf{Q}(x)$ has a finite dimension as a $\mathbf{Q}(x)$-vector space.
@@
has $p(x(t))\in\mathbf{Z}\llbracket t\rrbracket$, then the following dimension bound holds on $\mathcal{H}(U,x(t),\mathbf{Z})$ over $\mathbf{Q}(p(x))$:
(2.0.4) $\dim_{\mathbf{Q}(p(x))}\big(\mathcal{H}(U,x(t),\mathbf{Z})\otimes\mathbf{Q}(p(x))\big)\leq e\cdot\frac{\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}}{\log{|\varphi^{\prime}(0)|}},$
where $e=2.718\ldots$ is Euler’s constant.
@@
Our approach rests on the remark that $\varphi^{*}\mathcal{H}(U,x(t),\mathbf{Z})\subset\varphi^{*}\mathcal{H}(U,\mathbf{C})$ lies in the ring of formal power series fulfilling linear differential equations with analytic coefficients and no singularities on the closed disc $\overline{D(0,1)}$ and hence, by Cauchy’s theorem, it is contained by the ring $\mathcal{O}(\overline{D(0,1)})$ of holomorphic functions on $\overline{D(0,1)}$.
@@ 2.1. The auxiliary construction
We consider $f_{1},\ldots,f_{m}\in\mathbf{Q}\llbracket x\rrbracket$ an $m$-tuple of $\mathbf{Q}(p(x))$-linearly independent functions in the space $\mathcal{H}(U,x(t),\mathbf{Z})$, and proceed to establish the holonomy bound
(2.1.1) $m\leq e\cdot\frac{\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\textrm{Haar}}}{\log{|\varphi^{\prime}(0)|}}.$
@@
    all $a_{\mathbf{i,k}}\in\mathbf{Z}$ are rational integers bounded in absolute value by $\exp\big(\kappa C\alpha+o(\alpha)\big)$ for some constant $C\in\mathbf{R}$ depending only on $\max_{u\notin U}{1/|u|}$ and on the degree and height of the rational function $p(x)\in\mathbf{Q}(x)$.
@@
We expand our sought-for formal function in (2.1.3) into a formal power series in $\mathbf{Q}\llbracket\mathbf{x}\rrbracket$ and solve $\binom{\alpha+d}{d}\sim\alpha^{d}/d!$ linear equations in the $(mD)^{d}$ free parameters $a_{\mathbf{i,j}}$. By the formal inverse function expansion, the integrality condition $p(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ entails that $x(t)\in t+t^{2}\mathbf{Z}\llbracket t/M\rrbracket$ for some $M\in\mathbf{N}$ bounded in terms of the degree and height of the rational function $p(x)$. The result then follows from the classical Siegel lemma [BG06, Lemma 2.9.1], with the degree parameter choice
@@
2.2. Equidistribution
The key idea now is that upon substituting $x_{j}=\varphi(z_{j})$ into (2.1.3), the $d\to\infty$ equidistribution on the circle of the uniform independent and identically distributed points $z_{1},\ldots,z_{d}$ will normally get the constituent monomials in (2.1.3) to grow at most at the integrated exponential rate of $dD\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}$. It is this Monte Carlo (randomized numerical integration) principle that makes possible the integrated growth term — as opposed to the cruder and rather more straightforward supremum growth term, compare to [And89, VIII 1.6] — in our holonomy rank bound (2.0.4). Having $\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}$ instead of $\sup_{|z|=1}\log{|p\circ\varphi|}$ in (2.0.4) is a critical step in our proof of the unbounded denominators conjecture.
To carry out this program, we employ in § 2.4 André’s method [And89, And04] of extrapolating with the function

### v2

@@ 1.1.4. The dimension bound can be leveraged with growing level NN
The meaning of the requisite inflation rate is clarified in § 4, with Proposition 4.2.7 and Remark 4.2.10. It turns out that the logarithmically inflated holonomy rank (dimension) bound by $O(N^{3}\log{N})$ is sufficient for the desired proof by contradiction (but an $O(N^{3+1/\log{\log{N}}})$ or worse form of bound would not suffice); and this is what we ultimately prove. Getting to this degree of precision creates however some additional challenges. A straightforward elaboration of André’s original argument in [And89, Criterium VIII 1.6], taking the number of variables $d\to\infty$ and involving the $\sup_{|z|=r}\log{|\varphi|}$ growth term of loc.cit. in place of our mean (integrated) growth term $m(r,\varphi)$ in (2.0.7), leads quite easily to an $O(N^{5})$ dimension bound; and by a further work explicitly with the cusps of the Fuchsian uniformization $D(0,1)/\Gamma_{N}\cong\mathbf{C}\smallsetminus\mu_{N}$, and an appropriate Riemann map precomposition, it is possible to further reduce that down to an $O(N^{4})$. See Remark 5.2.18. This does not suffice to conclude the proof. Going further requires an intrinsic improvement into André’s dimension bound itself: the reduction of the supremum term to the integrated term in the numerator of (2.0.7).
We give three proofs of this improvement, all being based on the same auxiliary construction scheme of § 2.1. Our default treatment §§2.1, 2.2, 2.3 is based on Nevanlinna’s canonical factorization of meromorphic functions of bounded characteristic. Additionally, we also include in § 2.5 our original argument based on equidistribution ideas, and a simplified alternative path § 2.4 proposed to us by André and based on plurisubharmonicity and a lexicographic induction. The former variation has a potential-theoretic flavor familiar from the proof of Bilu’s theorem [Bil97] (see also § 2.5.4), but it is in the cross-variables $d\to\infty$ asymptotic aspect and hence different than the well-established link (see [Bos99, BCL09, Bos04]) of arithmetic algebraization to adelic potential theory.
@@ 2. The arithmetic holonomicity theorem
Our proof relies on the following dimension bound which is an extension of André’s arithmetic algebraicity criterion [And04, Théorème 5.4.3]. We state and prove our result here in a particular case suited to our needs, beginning with the abstract form. We denote by $\mathcal{O}(\overline{D(0,1}))\subset\mathbf{C}\llbracket z\rrbracket$ the ring of holomorphic function germs that converge on some open neighborhood of the closed unit disc $|z|\leq 1$. Throughout our paper, we will use the notation
@@
Consider the following data:
  • (i)
    a nonconstant rational function $p(x)\in\mathbf{Q}(x)\smallsetminus\mathbf{Q}$ without pole at $x=0$,
  • (ii)
    a formal power series
    (2.0.2) $x(t)\in t+t^{2}\mathbf{Q}\llbracket t\rrbracket$
    pulling back $p$ into an integral coefficients power series $x^{*}p:=p(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ in the new variable $t$,
  • (iii)
    and a holomorphic mapping $\varphi:\overline{D(0,1)}\to\mathbf{C}$ taking $\varphi(0)=0$ with $|\varphi^{\prime}(0)|>1$, and pulling back $p$ into a holomorphic function $\varphi^{*}p\in\mathcal{O}(\overline{D(0,1)})$ on some neighborhood of the closed unit disc.
Suppose the formal power series $f_{1},\ldots,f_{m}\in\mathbf{Q}\llbracket x\rrbracket$ are $\mathbf{Q}(p(x))$-linearly independent and satisfy the following integrality and analyticity properties like in (ii) and (iii):
$x^{*}f_{1},\ldots,x^{*}f_{m}\in\mathbf{Z}\llbracket t\rrbracket,\quad\textrm{and}\quad\varphi^{*}f_{1},\ldots,\varphi^{*}f_{m}\in\mathcal{O}(\overline{D(0,1)}).$
Then $f_{1},\ldots,f_{m}\in\mathbf{Q}\llbracket x\rrbracket$ are algebraic (i.e., all $f_{i}\in\overline{\mathbf{Q}(x)}$), and
(2.0.3) $m\leq e\cdot\frac{\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}}{\log{|\varphi^{\prime}(0)|}},$
where $e=2.718\ldots$ is Euler’s constant.
The novel point of the bound (2.0.3) is the integrated term in the numerator instead of a supremum term. This is critical for our proof of the unbounded denominators conjecture to have the numerator in (2.0.3), which measures the growth of $\varphi$, expressed as a Nevanlinna mean proximity function (or a characteristic function).
This abstract dimension bound (2.0.3) will be used more concretely as a holonomy rank bound. To state the relevant corollary, let us introduce an algebra of holonomic power series with integral coefficients and restricted singularities.
Definition 2.0.4.
For $U\subset\mathbf{C}$ an open subset, $R\subset\mathbf{C}$ a subring with fraction field $F:=\mathrm{Frac}(R)$, and $x(t)\in t\mathbf{Q}\llbracket t\rrbracket$ a formal power series, we define $\mathcal{H}(U,x(t),R)$ to be the ring of formal power series $f(x)\in F\llbracket x\rrbracket$ whose $t$-expansion $f(x(t))\in R\llbracket t\rrbracket$, and such that there exists a nonzero linear differential operator $L$ over $\overline{\mathbf{Q}}(x)$ with $L(f)=0$ and having a trivial local monodromy around all of its singular points that belong to $U$.
Further, we let $\mathcal{V}(U,x(t),R)$ to be the $F(x)$-vector space spanned by $\mathcal{H}(U,x(t),R)$.
For $x(t)=t$, we more simply denote the $R[x]$-algebra $\mathcal{H}(U,t,R)$ by $\mathcal{H}(U,R)$ and the $F(x)$-vector space $\mathcal{V}(U,t,R)$ by $\mathcal{V}(U,R)$.
Here by trivial local monodromy around $x=\alpha$ we mean that there exist a complex neighborhood $U_{\alpha}\ni\alpha$ and meromorphic functions $g_{1},\dots,g_{n}\in\mathcal{M}(U_{\alpha})$ on $U_{\alpha}$, where $n$ is the order of $L$, such that $g_{1},\dots,g_{n}$ form a $\mathbf{C}$-basis of the solution space of $L(f)=0$ on $U_{\alpha}\smallsetminus\{\alpha\}$. This is the case if $x=\alpha$ is not a singular point of $L$, and apart from that the typical examples singular at $x=0$ include $L_{n}=x\frac{d}{dx}-n$ for $n\in\mathbf{Z}\smallsetminus\{0\}$, of solution space $\ker{L_{n}}=\mathbf{C}\cdot x^{n}$.
Our holonomy bound is now a straightforward combination of Theorem 2.0.1 and Cauchy’s analyticity theorem on the solutions of linear differential equations with analytic coefficients.
Corollary 2.0.5.
Let $0\in U\subset\mathbf{C}$ be an open subset containing the origin. If the uniformization radius of the pointed Riemann surface $(U,0)$ is strictly greater than $1$, then the algebra $\mathcal{V}(U,\mathbf{Z})$ is finite dimensional as a $\mathbf{Q}(x)$-vector space.
@@
has $p(x(t))\in\mathbf{Z}\llbracket t\rrbracket$, then the following dimension bound holds on $\mathcal{V}(U,x(t),\mathbf{Z})$ over $\mathbf{Q}(p(x))$:
(2.0.7) $\dim_{\mathbf{Q}(p(x))}\mathcal{V}(U,x(t),\mathbf{Z})\leq e\cdot\frac{\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}}{\log{|\varphi^{\prime}(0)|}}.$
@@
The pulled back space $\varphi^{*}\mathcal{H}(U,x(t),\mathbf{Z})\subset\varphi^{*}\mathcal{H}(U,\mathbf{C})$ lies in the ring of formal power series fulfilling linear differential equations with analytic coefficients and no singularities with nontrivial local monodromies on the closed disc $\overline{D(0,1)}$. Hence, for any such function $f\in\mathcal{H}(U,x(t),\mathbf{Z})$, there exists a nonzero $g(x)\in\mathbf{Q}[p(x)]\smallsetminus\{0\}$ such that for any singular point $\alpha\in U$ of the linear operator $L$ in Definition 2.0.4, and for any local solution $h(x)$ in a small punctured neighborhood of $\alpha$, the product function $g(x)h(x)$ is holomorphic at $x=\alpha$. (The singularities of $L$ and thus $h(x)$ in $U$ are all at algebraic points, hence such a $g(x)$ exists.) Cauchy’s theorem then gives that $\varphi^{*}(gf)$ is a holomorphic function on $\overline{D(0,1)}$, and we conclude by Theorem 2.0.1. ∎
Theorem 2.0.1 is modeled on André’s Diophantine approximation method [And89, §VIII], [And04, §5]. We include as many as three proofs, all sharing a common basic framework § 2.1 and relying crucially on a $d\to\infty$ limit for the number of auxiliary variables in the auxiliary function constructed by Lemma 2.1.2 below. Our original treatment was based on equidistribution and is in §§2.1, 2.5, and an alternative approach proposed to us by André and based on plurisubharmonicity is in §§2.1, 2.4. Firstly we give a shorter proof based on Nevanlinna’s canonical factorization §2.3 and the following intermediate form of Theorem 2.0.1.
Theorem 2.0.8.
In the setting of Theorem 2.0.1, consider furthermore an arbitrary holomorphic function $h:\overline{D(0,1)}\to\mathbf{C}$ with $h(0)=1$. Then
(2.0.9) $m\leq e\,\frac{\max\big\{\sup_{\mathbf{T}}{\log{|h|}},\,\sup_{\mathbf{T}}{\log{|h\cdot\varphi^{*}p|}}\big\}}{\log{|\varphi^{\prime}(0)|}},$
and $f_{1},\ldots,f_{m}\in\overline{\mathbf{Q}(x)}\cap\mathbf{Q}\llbracket x\rrbracket$.
We invite the reader on a first pass to proceed directly to §3 after §2.3.
@@ 2.1. The auxiliary construction
Since $\varphi$ maps $(D(0,1),0)\to(\mathbf{C},0)$ with nonzero derivative, the inverse function theorem gives a positive radius $\rho>0$ such that
(2.1.1) $\varphi\,:\,\varphi^{-1}(D(0,\rho))_{0}\xrightarrow{\cong}D(0,\rho)$
is an analytic isomorphism from the connected component $\varphi^{-1}(D(0,\rho))_{0}$ containing 0 of $\varphi^{-1}(D(0,\rho))$.
@@
    all $a_{\mathbf{i,k}}\in\mathbf{Z}$ are rational integers bounded in absolute value by $\exp\big(\kappa C\alpha+o(\alpha)\big)$ for some constant $C\in\mathbf{R}$ depending only on the radius $\rho$ from (2.1.1) and on the degree and height of the rational function $p(x)\in\mathbf{Q}(x)$.
@@
We expand our sought-for formal function in (2.1.3) into a formal power series in $\mathbf{Q}\llbracket\mathbf{x}\rrbracket$ and solve $\binom{\alpha+d}{d}\sim\alpha^{d}/d!$ linear equations in the $(mD)^{d}$ free parameters $a_{\mathbf{i,j}}$. By the formal inverse function expansion, the integrality condition $p(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ entails that $x(t)\in t+(t^{2}/M)\mathbf{Z}\llbracket t/M\rrbracket$ for some $M\in\mathbf{N}$ bounded in terms of the degree and height of the rational function $p(x)$. Therefore also the inverse series $t(x)\in x+(x^{2}/M)\mathbf{Z}\llbracket x/M\rrbracket$, and so $f_{i}(x(t))\in\mathbf{Z}\llbracket t\rrbracket$ entails $f_{i}(x)\in\mathbf{Z}\llbracket x/M\rrbracket$ for all $i=1,\ldots,m$. Furthermore, by (2.1.1), every power series $f_{i}(x)\in\mathbf{Q}\llbracket x\rrbracket$ is convergent on the archimedean disc $|x|<\rho$. The result then follows from the classical Siegel lemma [BG06, Lemma 2.9.1], with $e^{C}:=M/\rho$ and the degree parameter choice
@@
2.2. Extrapolation and proof of Theorem 2.0.8
We consider the nonzero formal function